# Uniaxial Compression-Induced Anisotropy and Electronic Dimensionality in the Iron-Based Superconductor FeSe

### 01 Apr 2026 v1

@@ Abstract
The evolution of the superconducting transition temperature () in FeSe was investigated under in-plane, out-of-plane, and hydrostatic compression. For pressures up to 0.6 GPa, increases regardless of the compression mode, consistent with the suppression of nematic ordering. However, once nematicity is suppressed, exhibits a striking directional dependence: out-of-plane compression shows behavior similar to the hydrostatic case, with a sharp increase in , whereas in-plane compression suppresses superconductivity. First-principles calculations suggest that in-plane compression shifts a hybridized band of Se and Fe character so that it crosses the Fermi level along the -Z direction, leading to the emergence of an additional metallic band. This results in an increased three-dimensionality of the electronic structure and may be interpreted as a possible Lifshitz-type change in the Fermi surface. These findings indicate that the dimensionality of the electronic structure plays a key role in determining the response of FeSe under different compression modes.
@@
Figure 3(b) shows the pressure dependence of the tetragonal distortion obtained from first-principles calculations. To model the out-of-plane uniaxial strain, the parameter obtained from the ambient-pressure case was fixed during optimization. Similarly, was fixed for in-plane strain. Previous hydrostatic compression experiments showed that correlates with [27]. Here also, the behavior of correlates well with the dependence observed above 1 GPa in Fig. 3(a). This parameter can be regarded as an indicator of the structural dimensionality. A larger value of corresponds to a more three-dimensional structure, whereas a smaller value indicates a more two-dimensional character. This suggests that the structural dimensionality characterized by plays an important role in determining the superconducting behavior in the high-pressure regime. While predicts a satisfying trend for , the electronic band structure will ultimately govern the superconducting behavior.
Here, to investigate the pressure-mode dependence of the electronic structure, we performed a fat-band analysis based on maximally localized Wannier functions [42]. Figure 4(a–d) shows the orbitally decomposed band structures for each compression mode. Although the overall features of the band structures are very similar, a notable difference appears along the -Z direction near the Fermi level. In the case of the in-plane compression, shown in Fig. 4(c), a band crosses the Fermi level near the middle of the -Z path, whereas no such crossing is observed under the other compression modes. We find that this band has a hybridized character of Se and Fe orbitals. This indicates that the electronic structure under in-plane compression has a more three-dimensional character. In contrast, the other compression modes exhibit more two-dimensional electronic structures with flat bands along the -Z path. A significant discrepancy emerges between structural and electronic character. While in-plane compression results in a structurally more two-dimensional unit cell (smaller ), fat-band analysis confirms the emergence of a new metallic band along the -Z direction, rendering the electronic structure more three-dimensional.
@@
The compression-mode dependence of the low-energy electronic structure can be discussed as follows. The suppression of the nematic phase under pressure may enhance magnetic fluctuations, thereby promoting superconductivity. The additional metallic channel appearing under in-plane compression likely triggers a redistribution of orbital occupancy. This electronic reconstruction prevents the expected enhancement seen in hydrostatic conditions, suggesting that the preservation of electronic two-dimensionality is vital for high- superconductivity in Fe-chalcogenides. This behavior may also be interpreted as a possible Lifshitz-type reconstruction of the Fermi surface associated with the emergence of the additional metallic band.

### 03 Apr 2026 v2

@@ Abstract
The evolution of the superconducting transition temperature () in FeSe was investigated under in-plane, out-of-plane, and hydrostatic compression. For pressures up to 0.6 GPa, increases regardless of the compression mode, consistent with the suppression of nematic ordering. However, once nematicity is suppressed, exhibits a striking directional dependence: out-of-plane compression shows behavior similar to the hydrostatic case, with a sharp increase in , whereas in-plane compression suppresses superconductivity. First-principles calculations suggest that in-plane compression shifts a hybridized band of Se and Fe character so that it crosses the Fermi level along the -Z direction, leading to the emergence of an additional metallic band. This results in an increased three-dimensionality of the electronic structure and may be interpreted as a possible Lifshitz-type change in the Fermi surface.
@@
Figure 3(b) shows the pressure dependence of the tetragonal distortion obtained from first-principles calculations. To model the out-of-plane uniaxial strain, the parameter obtained from the ambient-pressure case was fixed during optimization. Similarly, was fixed for in-plane strain. Previous hydrostatic compression experiments showed that correlates with [27]. Here also, the behavior of correlates well with the dependence observed above 1 GPa in Fig. 3(a). This parameter can be regarded as an indicator of the structural dimensionality. A larger value of corresponds to a more three-dimensional structure, whereas a smaller value indicates a more two-dimensional character. This suggests that the structural dimensionality characterized by plays an important role in determining the superconducting behavior in the high-pressure regime. While predicts a consistent trend for , the electronic band structure will ultimately govern the superconducting state.
Here, to investigate the pressure-mode dependence of the electronic structure, we performed a fat-band analysis based on maximally localized Wannier functions [42]. Figure 4(a–d) shows the orbitally decomposed band structures for each compression mode. Although the overall features of the band structures are very similar, a notable difference appears along the -Z direction near the Fermi level. In the case of the in-plane compression, shown in Fig. 4(c), a band crosses the Fermi level near the middle of the -Z path, whereas no such crossing is observed under the other compression modes. This band is found to have a hybridized character of Se and Fe orbitals, as shown in Figure 4(a–d). This indicates that the electronic structure under in-plane compression has a more three-dimensional character. In contrast, the other compression modes exhibit more two-dimensional electronic structures with flat bands along the -Z path. A significant discrepancy emerges between structural and electronic character. While in-plane compression results in a structurally more two-dimensional unit cell (smaller ), fat-band analysis confirms the emergence of a new metallic band along the -Z direction, rendering the electronic structure more three-dimensional.
@@
The compression-mode dependence of the low-energy electronic structure can be discussed as follows. The suppression of nematic ordering under pressure may enhance magnetic fluctuations, which could in turn promote superconductivity. The additional metallic channel appearing under in-plane compression likely triggers a redistribution of orbital occupancy. This electronic reconstruction prevents the expected enhancement seen in hydrostatic conditions, suggesting that the preservation of electronic two-dimensionality is vital for high- superconductivity in Fe-chalcogenides. This behavior may also be interpreted as a possible Lifshitz-type reconstruction of the Fermi surface associated with the emergence of the additional metallic band.